the density of Cu is 8.94 g/cm3. The quantity of electricity needed to plate an area 100cm2 to a thikness of 1/102 cm using  CuSo4 solution is ?

Grade:12th pass

2 Answers

Abhishek Singh
9 years ago
vol. of Cu to be plated =100*1/10=10cc
wt. of cu reqd.= 8.94*10=89.4
moles of cu = 89.4/63.5
Cu2++2e-----------.> Cu
Moles of electron = 89.4*10*2
amt. of electricity= 89.4*10*2/96500 Farads
Pravin Sudhakar Mahajan
7 years ago
Volume (V) of Cu = Area x thickness = 10x10 x 1/102 = 1.
d = m/V
m = d x V= 8.94 x 1 = 8.94.
Hence, Actual mass of Cu to be plated is 8.94 g.
Now, Cu2+ + 2e- -------> Cu
So, 2F --------> 63.5 g Cu
       2 x 96500 C -------> 63.5g Cu
                   P  C -------> 8.94 g Cu
                   P = 2 x 96500 x 8.94/96500 = 27172.
Hence, the answer is 27172